OpenStudy (anonymous):

Solve 5 log x = -15? I'm really stuck. :/

OpenStudy (anonymous):

Not sure if this is right... try doing it on mathway.com

OpenStudy (lgbasallote):

hint: divide both sides by 5 first. what does that give you?

OpenStudy (anonymous):

Would it be -3? :/

OpenStudy (lgbasallote):

yes. \[\log x = -3\] now change this into exponential form. do you know how?

OpenStudy (anonymous):

No, that's where I'm really getting stumped. :/ Thanks for helping me. This is great.

OpenStudy (lgbasallote):

here's an example..when you have \[\log 2 = x\] this becomes \[10^x = 2\] because if the base of the log is not given, it's automatically 10. do you get what to do now?

OpenStudy (anonymous):

Umm.. so would I multiply -3 by 10 then? Or am I getting that wrong?

OpenStudy (lgbasallote):

no.. you raise 10 to -3 \[\LARGE 10^{-3} = x\]

OpenStudy (lgbasallote):

so what is the value of 10^-3?

OpenStudy (anonymous):

Would it be 1000? :)

OpenStudy (lgbasallote):

hmm close but no.. 1000 is 10^3 \[\Large 10^{-3} \implies \frac{1}{10^3}\] would you like to try again?

OpenStudy (anonymous):

Oh okay. So then would it be .001?

OpenStudy (lgbasallote):

right

OpenStudy (lgbasallote):

or simply 1/1000

OpenStudy (lgbasallote):

whatever floats the boat

OpenStudy (anonymous):

Great!!! Thank you so much. I really appreciate your help.

OpenStudy (lgbasallote):

welcome